OpenStudy (anonymous):

Balance the combustion reaction between octane and oxygen. 2C^8H^8+___O^2→____CO^2+___H^2O

OpenStudy (ash2326):

@mapa_ramos This section is for maths questions, next time please post Chemistry Questions in Chemistry:) Thanks Nevertheless I'll help you here!!

OpenStudy (ash2326):

\[2C_8H_8+O_2 \longrightarrow CO_2+H_2O\] Tell me how many carbon atoms are on the left side?

OpenStudy (anonymous):

I know, but in the chemistry Questions section nobody helps me...umm 2?

OpenStudy (ash2326):

\[2C_8=> 2\times 8=16\], so I require 16 C on the right side also, I multiply 16 to C \[2C_8H_8+O_2 \longrightarrow \overbrace {16}CO_2+H_2O\] Do you get this?

OpenStudy (anonymous):

yes, I did get that

OpenStudy (ash2326):

Now tell me the no. of Hydrogen on left and right side?

OpenStudy (anonymous):

So hydrogen would be 2*9=18 so on the other side it would be 9H^2?

OpenStudy (ash2326):

It's 2*8 on the left side

OpenStudy (anonymous):

yeah, but on the other side left side second space..ok I`m lost again

OpenStudy (ash2326):

16 on the left side, right side we have 2, so we multiply by 8 (16=8*2) \[2C_8H_8+O_2 \longrightarrow {16}CO_2+\overbrace{8}H_2O\]

OpenStudy (ash2326):

Now you find the no. of Oxygen on the left side and right side?

OpenStudy (anonymous):

so there is O^2 on the left side and O o the right side?

OpenStudy (anonymous):

*on

OpenStudy (ash2326):

\[2C_8H_8+O_2 \longrightarrow {16}CO_2+{8}H_2O\] 2 O on left and (16*2+8=40) on the right so what should we multiply left side \(O_2\) with??

OpenStudy (anonymous):

20?

OpenStudy (ash2326):

Yeah you're right, so our balanced equation is \[ \huge\color {gold }{ {2C_8H_8+20O_2 \longrightarrow {16}CO_2+8H_2O}}\]

OpenStudy (anonymous):

okay I got everything except to the 8H^2O, how did you get that one since the other side is just H^2?

OpenStudy (ash2326):

Left side \[2 \times 8=16\] Right side we have \(H_2\) so 2 we need 16 both sides so I multiply 8

OpenStudy (anonymous):

ohh okay, thank you very much :)

OpenStudy (ash2326):

did you understand?

OpenStudy (anonymous):

yes :)